OpenStudy (darleenie):

x^2-7x+12=0 complete the square

OpenStudy (holsteremission):

Hint: \(x^2-7x+\dfrac{49}{4}=\left(x-\dfrac{7}{2}\right)^2\) What can you do to the left hand side of the equation to isolate a factor of \(\dfrac{49}{4}\)?